40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat May 9 00:15:21 2009, Yoichi, Update, PSL, Laser head temperature oscillation laser.png
    Reply  Sat May 9 02:20:11 2009, Jenne, Update, PSL, Laser head temperature oscillation 
       Reply  Sat May 9 15:19:10 2009, rana, Update, PSL, Laser head temperature oscillation Picture_2.png
          Reply  Mon May 11 11:49:20 2009, steve, Update, PSL, MOPA cooling water lines are backwards htempvar80d.jpg
             Reply  Thu May 21 18:07:32 2009, rana, Update, PSL, Screw on Needle valve loosened Untitled.png
Message ID: 1568     Entry time: Sat May 9 00:15:21 2009     Reply to this: 1569
Author: Yoichi 
Type: Update 
Category: PSL 
Subject: Laser head temperature oscillation 
After the laser cooling pipe was unclogged, the laser head temperature has been oscillating in 24h period.
The laser power shows the same oscillation.
Moreover, there is a trend that the temperature is slowly creeping up.
We have to do something to stop this.
Or Rob has to finish his measurements before the laser dies.
Attachment 1: laser.png  154 kB  | Hide | Hide all
laser.png
ELOG V3.1.3-